OpenStudy (anonymous):

Find the standard form of the equation of the hyperbola with the given characteristics. vertices: (2, - 6), (2, - 2) foci: (2, - 8), (2, 0) I know this much so far: (y+4)^2 (x-2)^2 ------ - ------ ? ?

OpenStudy (amistre64):

the y is the heavy side; so this opens up and down

OpenStudy (amistre64):

teh distance between the foci is 8; so each foci is 4 from center; that 4 is either a or c :)

OpenStudy (anonymous):

ok so a

OpenStudy (amistre64):

we have 4 between the vertices; so each vertex is 2 from center

OpenStudy (anonymous):

right how do I find b^2 then?

OpenStudy (amistre64):

under the y we have 2^2 right? let me look this up some more :)

OpenStudy (anonymous):

ok I think actually I have to know what c^2 is cause C^2=a^2+b^2

OpenStudy (anonymous):

I'm thinkin' b^2=16...?

OpenStudy (amistre64):

tell me your understanding of a b and c; what do they measure?

OpenStudy (anonymous):

the foci are c units from the center the vertices are a units from the center

OpenStudy (amistre64):

so; a^2 + b^2 = c^2 then right?

OpenStudy (amistre64):

2 + b^2 = 64 b^2 = 62 then

OpenStudy (anonymous):

I think so....but that doesn't eem right maybe it's minus?

OpenStudy (anonymous):

my options for b^2 are 12 or 16 here but I don't see how they get either of those

OpenStudy (amistre64):

like this right?

OpenStudy (amistre64):

i used 8 instead of 4 :)

OpenStudy (anonymous):

i think so...

OpenStudy (amistre64):

16-4 = 12

OpenStudy (amistre64):

b^2 = 12

OpenStudy (anonymous):

oh ok I was gonna say 16 cause 4^2=16 but 12 works too

OpenStudy (amistre64):

the minor axis is always smaller in number than the major axis

OpenStudy (amistre64):

since the major is the y on this one; and it is 16; the other has to be less

OpenStudy (amistre64):

or maybe i messed that up in thougth

OpenStudy (anonymous):

ugh... idk this is all a review for my final and I am thinking that so far i'm not doing too hot. YIKES

OpenStudy (amistre64):

y^2 x^2 --- - --- = 1 4 12 b^2 = focal distance(c)^2 - vertex distance(a)^2 in this case; the hyperbola; c is the greatest distance beacusae it is used as the hypotenust in the ellipse; a is the greatest distance becasue IT is used for ITS hypotenuse